# Rydberg superatoms: An artificial quantum system for quantum information processing and quantum optics

### v1

@@ V.3 Hybrid Rydberg system
Hybrid systems involving Rydberg atoms promise intriguing possibilities for quantum information processing. It was shown that a Jaynes-Cummings (JC) model could be realized in the optical domain for an ensemble of atoms by exploiting their collective couplings to a quantized cavity mode.Guerlin et al. (2010); Shao, Wu, and Yi (2017) This is realized by restricting the huge state space to the collective ground state $|G\rangle$ and the collective excited state $|E\rangle$ with a single Rydberg excitation due to dipole blockade. The main benefit is that the collectively enhanced coupling can be much larger than the decay rates, leading to the strong coupling regime of an effective JC model. Trapping some atoms in a volume smaller than the blockade sphere close to an atom chip, it is viable to attain a hybrid system consisting of a superatom coupled to a surface phonon polariton, which can be used to subtract individual photons from a beam of light.Kübler et al. (2013) Hybrid systems consisting of macroscopic mechanical oscillators coupled to a small ensemble of Rydberg atoms are also widely investigated to achieve ground state cooling, quantum state engineering, nontrivial nonlinear dynamics, bistable Rydberg excitation, and mechanical oscillation, etc.Yan et al. (2015); Bariani et al. (2014); Carmele et al. (2014) This depends especially on the collectively enhanced coupling arising from the dipole blockade effect of Rydberg superatoms.Stiesdal et al. (2018); Yang et al. (2022b) The strong coupling regime may also be realized between a Rydberg atomic ensemble and propagating surface phonon polaritons on a piezoelectric superlattice.Sheng, Chao, and Shaffer (2016) These hybrid Rydberg systems provide a promising platform for extending the contents of nonlinear quantum optics and photonics, essential for developing future quantum control techniques. Moreover, when combined with Floquet engineering techniques, polaritons, formed by photons entering a cavity and hybridized with atomic Rydberg excitations, can be efficiently manipulated to interact with each other.Clark et al. (2019) They can also be simulated to exhibit topological order, such as the Laughlin state.Clark et al. (2020)
@@
Figure 27: The experimental setup involves measuring transmission, with EIT spectra depicted in blue (presence of EIT control) or red (absence of EIT control), normalized against the maximum transmission of an empty cavity (depicted in gray). Reprinted with permission from Vaneecloo et al., Phys. Rev. X 12(2), 021034 (2022). Copyright 2022 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Vaneecloo, Garcia, and Ourjoumtsev (2022)
@@
Figure 28: (a) A bow-tie cavity composed of two concave mirrors and two convex mirrors. Relative intensity (b) and (c) phase of signal light as a function of signal detuning referring to cavity-enhanced Rydberg EIT spectra. Red (blue) data and lines are attained in the absence (presence) of coupling light. Reprinted with permission from Stolz et al., Phys. Rev. X 12(2), 021035 (2022). Copyright 2022 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Stolz et al. (2022)
@@
Below, we will explain the above discussions in more detail by focusing on recent experiments.Vaneecloo, Garcia, and Ourjoumtsev (2022); Stolz et al. (2022); Magro et al. (2023) As can be seen in Fig. 27, by placing a small cloud of ${}^{87}$b atoms inside a cavity of finesse $\mathcal{F}=590$, one observes the vacuum Rabi splitting of two transmission peaks located at $\delta/2\pi=\pm 10$ MHz, indicating a cavity-mode coupling constant ($g/2\pi=10$ MHz) exceeding both cavity decay rate ($\kappa/2\pi=2.9$ MHz) and atomic decay rate ($\gamma/2\pi=3$ MHz). Exciting these atoms to a high Rydberg state $|R^{\prime}\rangle$ by the EIT control and D1 probe beams, a narrower EIT window is found to open on resonance corresponding to a dark polariton state mixing a photon in the cavity and a Rydberg excitation in the atomic cloud. On the other hand, the $109S$ drive and D2 drive beams can drive these atoms to another high Rydberg state $|R\rangle$ via a two-photon Raman resonance so that they behave as a two-level superatom due to the blockade of multiple Rydberg excitations. Atoms in states $|R\rangle$ and $|R^{\prime}\rangle$ further strongly interact via their dipoles as a result of the Förster resonance with another pair of Rydberg states. Then, by shifting the state $|R\rangle$ out of resonance, a superatom in the state $|R^{\prime}\rangle$ destroys the EIT response to the D1 probe, which drives the photonic component of the $|R^{\prime}\rangle$ polariton. This scheme has been exploited to realize the high-efficiency single shot optical detection of a Rydberg superatom, the superatom state-dependent $\pi$ phase shift of light reflected by the cavity, and the deterministic preparation of freely propagating photonic qubits with negative Wigner functions,Vaneecloo, Garcia, and Ourjoumtsev (2022); Magro et al. (2023) by taking advantage of two complementary approaches: i.e., cavity quantum electrodynamics and interacting atomic ensembles.Kumar et al. (2016)
@@
It is also worth noting that a bow-tie cavity has been incorporated into a Rydberg EIT experiment for demonstrating a polarization CNOT gate between two optical pulses interacting with hundreds of trapped ${}^{87}$b atoms at an average efficiency of about $41.7\%$, far exceeding that in previous experiments.Stolz et al. (2022) This atom-cavity system in a ring geometry minimizes dephasings resulting from the combination photon recoil of atomic motion and is characterized by $(g,\kappa,\gamma)=2\pi\times(1.0,2.3,3.0)$ MHz, thus yielding finesse $\mathcal{F}=350$ and a collective cooperativity of $21$ facilitating the enhanced interaction of single photons and Rydberg atoms. The four-mirror bow-tie cavity is shown in Fig. 28(a) where the target and control coupling beams are guided into the cavity from different concave mirrors, while the target and control qubits are input from a common convex mirror. This ring resonator is necessary for converting the $\pi$ conditional phase shift into a two-photon gate for polarization qubits. The relative intensity and phase of the target signal reflected from the cavity have been measured as a function of the signal detuning from the cavity resonance in Figs. 28(b) and 28(c), respectively. In the absence of an EIT coupling light, the intensity exhibits a normal-mode splitting, while the phase shows two dispersive features, one at a normal mode (see red data). As the EIT coupling light is on, a narrow EIT feature appears at resonance so that a $2\pi$ phase boost may occur as the signal detuning is tuned through zero (see blue data) - the origin of a $\pi$ conditional phase shift on which the CNOT gate could be built.
@@ VI CONCLUSION AND OUTLOOK
In conclusion, Rydberg atoms have emerged as one of the most promising platforms for probing quantum physics and quantum technological applications. The essential elements necessary for understanding Rydberg physics are covered in the first part of this review article, particularly the optical excitation of Rydberg states and the emergence of different types of Rydberg-Rydberg interactions, such as van der Waals and dipolar interactions, including the tunability using external fields. The strong interatomic interactions lead to the emergence of superatoms in Rydberg gases via the Rydberg blockade. When driving by resonant lasers, many-body Rabi oscillations can be found. The superatom approach is an efficient way to describe the dynamics and collective states, while the underlying mechanism can be used to generate robust and scalable entanglement in neutral atoms. Further, the same mechanism can create the entanglement between atoms and photons and between two photons. In an atom array, Rydberg blockade and superatoms serve a mechanism to emulate many-body spin models to create various many-body states and intricate quantum dynamics.
@@ ACKNOWLEDGMENTS
The authors would like to extend their sincere gratitude to F. Cesa, C. S. Adams, S. Dürr, S. Hofferberth, Y. O. Dudin, and D. Petrosyan, for their valuable comments. This work was supported by the National Natural Science Foundation of China (NSFC) under Grant Nos. 12174048, 12274376, 12074061, 12374329, and U21A6006, and the National Key Research and Development Program of China under Grant No. 2021YFA1402003. R.N. acknowledges DST-SERB for Swarnajayanti fellowship File No. SB/SJF/2020-21/19, and the MATRICS grant (MTR/2022/000454) from SERB, Government of India. W.L. acknowledges support from the EPSRC through Grant No. EP/W015641/1 and the Going Global Partnerships Programme of the British Council (Contract No. IND/CONT/G/22-23/26).
@@
Xiao-Qiang Shao: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Shi-Lei Su: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Lin Li: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Rejish Nath: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Jin-Hui Wu: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Weibin Li: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal).

### v2

@@ V.3 Hybrid Rydberg systems
Hybrid quantum systems are systems that integrate various types of quantum components or platforms to capitalize on their complimentary qualities. These systems seek to exploit the distinct qualities of diverse quantum technologies in order to produce more robust, scalable, and adaptable quantum devices. Hybrid systems involving Rydberg atoms promise intriguing possibilities for quantum information processing. It was shown that a JC model could be realized in the optical domain for an ensemble of atoms by exploiting their collective couplings to a quantized cavity mode.Guerlin et al. (2010); Shao, Wu, and Yi (2017) This is realized by restricting the huge state space to the ground state $|G\rangle$ and the collective excited state $|R\rangle$ with a single Rydberg excitation due to Rydberg blockade. The main benefit is that the collectively enhanced coupling between the emitter and the cavity mode can be much larger than the decay rates, leading to the strong coupling regime of an effective JC model. Trapping some atoms in a volume smaller than the blockade sphere close to an atom chip, it is viable to attain a hybrid system consisting of a superatom coupled to a surface phonon polariton, which can be used to subtract individual photons from a beam of light.Kübler et al. (2013) Hybrid systems consisting of macroscopic mechanical oscillators coupled to a small ensemble of Rydberg atoms are also widely investigated to achieve ground state cooling, quantum state engineering, nontrivial nonlinear dynamics, bistable Rydberg excitation, and mechanical oscillation, etc.Yan et al. (2015); Bariani et al. (2014); Carmele et al. (2014) This depends especially on the collectively enhanced coupling arising from the dipole blockade effect of Rydberg superatoms.Stiesdal et al. (2018); Yang et al. (2022b) The strong coupling regime may also be realized between a Rydberg atomic ensemble and propagating surface phonon polaritons on a piezoelectric superlattice.Sheng, Chao, and Shaffer (2016) These hybrid Rydberg systems provide a promising platform for extending the contents of nonlinear quantum optics and photonics, essential for developing future quantum control techniques. Moreover, when combined with Floquet engineering techniques, polaritons, formed by photons entering a cavity and hybridized with atomic Rydberg excitations, can be efficiently manipulated to interact with each other.Clark et al. (2019) They can also be simulated to exhibit topological order, such as the Laughlin state.Clark et al. (2020)
@@
Figure 27: The experimental setup involves measuring transmission, with EIT spectra depicted in blue (presence of EIT control) or red (absence of EIT control), normalized against the maximum transmission of an empty cavity (depicted in gray). Reproduced with permission from Vaneecloo et al., Phys. Rev. X 12(2), 021034 (2022). Copyright 2022 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Vaneecloo, Garcia, and Ourjoumtsev (2022)
@@
Figure 28: (a) A bow-tie cavity composed of two concave mirrors and two convex mirrors. Relative intensity (b) and (c) phase of signal light as a function of signal detuning referring to cavity-enhanced Rydberg EIT spectra. Red (blue) data and lines are attained in the absence (presence) of coupling light. Reproduced with permission from Stolz et al., Phys. Rev. X 12(2), 021035 (2022). Copyright 2022 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Stolz et al. (2022)
@@
Below, we will explain the above discussions in more detail by focusing on recent experiments.Vaneecloo, Garcia, and Ourjoumtsev (2022); Stolz et al. (2022); Magro et al. (2023) As can be seen in Fig. 27, by placing a small cloud of 87Rb atoms inside a cavity with a finesse of $\mathcal{F}=590$, one observes the vacuum Rabi splitting of two transmission peaks located at $\delta/2\pi=\pm 10$ MHz, indicating a cavity-mode coupling constant ($g/2\pi=10$ MHz) exceeding both the cavity decay rate ($\kappa/2\pi=2.9$ MHz) and the atomic decay rate ($\gamma/2\pi=3$ MHz). Exciting these atoms to a high Rydberg state $|R^{\prime}\rangle$ by the EIT control and D1 probe beams, a narrower EIT window is found to open on resonance corresponding to a dark polariton state mixing a photon in the cavity and a Rydberg excitation in the atomic cloud. On the other hand, the $109S$ drive and D2 drive beams can drive these atoms to another high Rydberg state $|R\rangle$ via a two-photon Raman resonance so that they behave as a two-level superatom due to the blockade of multiple Rydberg excitations. Atoms in states $|R\rangle$ and $|R^{\prime}\rangle$ further strongly interact with another pair of Rydberg states through Förster resonance. Then, by shifting the state $|R\rangle$ out of resonance, a superatom in the state $|R^{\prime}\rangle$ destroys the EIT response to the D1 probe, which drives the photonic component of the $|R^{\prime}\rangle$ polariton. This scheme has been exploited to realize the high-efficiency single shot optical detection of a Rydberg superatom, the superatom state-dependent $\pi$ phase shift of light reflected by the cavity, and the deterministic preparation of freely propagating photonic qubits with negative Wigner functions,Vaneecloo, Garcia, and Ourjoumtsev (2022); Magro et al. (2023) by taking advantage of two complementary approaches: i.e., cavity quantum electrodynamics and interacting atomic ensembles.Kumar et al. (2016)
@@
It is also worth noting that a bow-tie cavity has been incorporated into a Rydberg EIT experiment for demonstrating a polarization CNOT gate between two optical pulses interacting with hundreds of trapped 87Rb atoms at an average efficiency of about $41.7\%$, far exceeding that in previous experiments.Stolz et al. (2022) This atom-cavity system in a ring geometry minimizes dephasing resulting from the combination photon recoil of atomic motion and is characterized by $(g,\kappa,\gamma)=2\pi\times(1.0,2.3,3.0)$ MHz, thus yielding finesse $\mathcal{F}=350$ and a collective cooperativity of $21$ facilitating the enhanced interaction of single photons and Rydberg atoms. The four-mirror bow-tie cavity is shown in Fig. 28(a) where the target and control coupling beams are guided into the cavity from different concave mirrors, while the target and control qubits are input from a common convex mirror. This ring resonator is necessary for converting the $\pi$ conditional phase shift into a two-photon gate for polarization qubits. The relative intensity and phase of the target signal reflected from the cavity have been measured as a function of the signal detuning from the cavity resonance in Figs. 28(b) and 28(c), respectively. In the absence of an EIT coupling light, the intensity exhibits a normal-mode splitting, while the phase shows two dispersive features, one at a normal mode (see red data). As the EIT coupling light is on, a narrow EIT feature appears at resonance so that a $2\pi$ phase boost may occur as the signal detuning is tuned through zero (see blue data). This forms the basis for a $\pi$ conditional phase shift, essential for constructing a CNOT gate.
@@ VI CONCLUSION AND OUTLOOK
In conclusion, Rydberg atoms have emerged as one of the most promising platforms for probing quantum physics and quantum technological applications. The essential elements necessary for understanding Rydberg physics are covered in the first part of this review article, particularly the optical excitation of Rydberg states and the emergence of different types of Rydberg-Rydberg interactions, such as vdw and dipolar interactions, including the tunability using external fields. The strong interatomic interactions lead to the emergence of superatoms in Rydberg gases via the Rydberg blockade. When driving by resonant lasers, many-body Rabi oscillations can be found. The superatom approach is an efficient way to describe the dynamics and collective states, while the underlying mechanism can be used to generate robust and scalable entanglement in neutral atoms. Further, the same mechanism can create the entanglement between atoms and photons and between two photons. In an atom array, Rydberg blockade and superatoms serve a mechanism to emulate many-body spin models to create various many-body states and intricate quantum dynamics.
@@ ACKNOWLEDGMENTS
The authors extend their sincere gratitude to the anonymous reviewers for their thorough reading and valuable suggestions, which have significantly improved the readability of the manuscript. Additionally, we would like to thank F. Cesa, C. S. Adams, S. Dürr, S. Hofferberth, Y. O. Dudin, D. Petrosyan, J. Zeiher, and A. Ourjoumtsev for their valuable comments. This work was supported by the National Natural Science Foundation of China (NSFC) under Grant Nos. 12174048, 12274376, 12074061, 12374329, and U21A6006, and the National Key Research and Development Program of China under Grant No. 2021YFA1402003. R.N. acknowledges DST-SERB for Swarnajayanti fellowship File No. SB/SJF/2020-21/19, the MATRICS grant (MTR/2022/000454) from SERB, and funding from National Mission on Interdisciplinary Cyber-Physical Systems (NM-ICPS) of the Department of Science and Technology, Govt. of India through the I-HUB Quantum Technology Foundation, Pune INDIA. W.L. acknowledges support from the EPSRC through Grant No. EP/W015641/1 and the Going Global Partnerships Programme of the British Council (Contract No. IND/CONT/G/22-23/26).
@@
Xiao-Qiang Shao: Project administration (equal); Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Shi-Lei Su: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Lin Li: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Rejish Nath: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Jin-Hui Wu: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Weibin Li: Project administration (equal); Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal).